OpenStudy (anonymous):

Determine the pH of a .19 m solution of pyridinium nitrate (C5H5NHNO3) at 25oC. [Pyridinium nitrate dissociates in water to give pyridinium ions (C5H5NH+), the conjugate acid of pyridine (Kb= 1.7x10^-9), and nitrate ions (NO3-).]

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@Photon336

OpenStudy (anonymous):

@Cuanchi

OpenStudy (cuanchi):

this looks like a weak base in solution will produce OH- calculate \[[OH-]= \sqrt{K _{b} \times C _{b} } \]

OpenStudy (anonymous):

\[\sqrt{1.7*10^{-9} *19}\]

OpenStudy (anonymous):

So something like that?

OpenStudy (cuanchi):

be carefull!! 0.19 M not 19 M

OpenStudy (anonymous):

Ooops sorry. I meant .19

OpenStudy (anonymous):

\[\sqrt{1.7*10^{-9} * .19}\]

OpenStudy (cuanchi):

yea then you calculate the log and subtract from 14 to get the pH

OpenStudy (anonymous):

=1.797x10^-5

OpenStudy (anonymous):

for the log does it need ot be -log or just log?

OpenStudy (cuanchi):

- log [OH-]= pOH pH+pOH= 14

OpenStudy (anonymous):

pH =-9.25

OpenStudy (anonymous):

Does that seem right?

OpenStudy (cuanchi):

no, the pH has to be positive!!!!

OpenStudy (anonymous):

So 9.25

OpenStudy (cuanchi):

sounds fine to me!

OpenStudy (anonymous):

Thanks